OpenStudy (anonymous):

Given the following system of equations, identify the type of system. x + y = 6 y = 3 - x consistent inconsistent equivalent

imqwerty (imqwerty):

do u know the what is a consistent and an inconsistent system?

Directrix (directrix):

This chart may help. @hhelpplzzzz

OpenStudy (pawanyadav):

1st is x+y=6 2nd is x+y=3 a1/a2=b1/b2 don't equal c1/c2c2 Where a,b are coefficient of x,y and c is constant of corresponding equation. So it is inconsistent.